Calculate the force of gravity between a newborn baby of mass 4 kg and the obstetrician of mass 70 kg, who is 0.2 m from the baby.
1
Expert's answer
2013-02-25T11:26:46-0500
Force of gravity, bydefinition, is expressed as
F = G*M*m/r^2 = 6.67*10^(-11) m^3/(kg*s^2) * 70 kg * 4 kg / (0.2 m)^2 = 4.669 *
10^(-7) N

Answer: approx. 467 nN (nanonewtons)